OpenStudy (anonymous):

Please Help! Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (mathstudent55):

Were you asked to use any specific method?

OpenStudy (solomonzelman):

subtract them from each other (elimination)

OpenStudy (anonymous):

No I was not. @mathstudent55

OpenStudy (anonymous):

Go to Mathway.com

OpenStudy (solomonzelman):

I can show you how to solve for one variable, and you will do the other , OK?

OpenStudy (solomonzelman):

-2x + y = 1 -4x + y = -1 ------------------- subtracting the 2nd equation from the first -2x-(-4x) +y-y = 1-(-1) -2x+4x + 0y = 1+1 -2x=2 x=-1 can you do the y?

OpenStudy (anonymous):

Yes I can! Thank you!(:

OpenStudy (solomonzelman):

Anytime, do you get the idea of how to do it in the future?

OpenStudy (anonymous):

Yes, Yes I doo.

OpenStudy (mathstudent55):

If not, then you can try elimination. Multiply the first equation by -1, and add it to the second equation. That will eliminate the y variable. Solve for x. Then use that value of x in one of the original equations to solve for y.